# Locking-Free Training of Physics-Informed Neural Network for Solving Nearly Incompressible Elasticity Equations

### v1

@@ 4 Numerical Experiments
Figure 4: A comparison of solutions in two dimensions when $\lambda=10^{5}$: the exact solution (a), the solution computed by the standard PINN (b), and the solution obtained by our method (c).
@@
Figure 5: A comparison of learning curves with respect to the relative $L^{2}$-error between the standard PINN and our method for various values of $\lambda$ in two dimensions.
In this section, we conduct a series of numerical experiments to assess the effectiveness of the proposed method. We compare the model trained using the standard PINN with the model trained with the proposed decomposition technique. To provide a comprehensive evaluation of the proposed method when the Lamé parameters $\mu$ and $\lambda$ in the elasticity equations are constant, including scenarios with large $\lambda$ (nearly incompressible material), we perform some numerical experiments in both two and three dimensions. We then investigate the case of variable Lamé coefficients $\lambda$ and $\mu$ (inhomogeneous material). Finally, we will also conduct experiments on a parametric PINN, which enables the rapid prediction of solutions for varying Lamé parameters without the need for retraining. To ensure a fair comparison, all other settings, including the model architecture, initialization, and optimization algorithm, remain unchanged.
@@ 4.1 Example 1 (Two-dimensional case)
${\bf u}(\boldsymbol{x})=\begin{bmatrix}u_{1}(x,y)\\ u_{2}(x,y)\end{bmatrix}=\begin{bmatrix}\bigl{(}\cos(2x)-1\bigr{)}\sin(2y)\\ \bigl{(}1-\cos(2y)\bigr{)}\sin(2x)\end{bmatrix}+\frac{\sin(x)\sin(y)}{\lambda}\begin{bmatrix}1\\ 1\end{bmatrix},$
where ${\boldsymbol{x}}=(x,y)\in\Omega$. For a baseline model, we implement a neural network consisting of four hidden layers, each containing 50 neurons, activated by the GELU function. For training, as described before, we adopt the decomposed loss function (3.5) with weights $\delta_{r}=0.05$, $\delta_{s}=1$ and $\delta_{t}=20$, and number of sampling points $N_{r}=5000$, $N_{s}=5000$ and $N_{t}=400$. We use the Adam optimizer as an optimization algorithm.
Figure 4 illustrates the solutions predicted by a standard PINN and our proposed method when $\lambda=10^{5}$ (nearly incompressible material). As shown in the figure, the solution obtained using the original PINN is close to zero with $1.01$e00 relative $L^{2}$-error, failing to accurately predict the solution of the given problem. In contrast, the solution computed using our proposed method closely approximates the exact solution with $1.65$e33-3- 3 relative $L^{2}$-error. To compare the learning curves for different values of $\lambda$, we conducted training using both the standard PINN and our proposed method for various values of $\lambda$, computing the relative $L^{2}$-error at each training epoch. The results of this experiment are presented in Figure 5. As observed from Figure 5 (a), when $\lambda=10$ (no locking is anticipated), there is no significant difference in training efficiency between the original PINN and our proposed method. However, as $\lambda$ is getting large, it is clear from (b) and (c) of Figures 5 that our method demonstrates superior training efficiency and accuracy compared to the standard PINN. It is noteworthy that, as we can see from Figure 5 (c) (when $\lambda=10^{5}$, the material becomes nearly incompressible), the standard PINN fails to train due to the non-robustness phenomenon in the nearly incompressible case, whereas our proposed method continues to achieve effective training with high accuracy.
@@
Figure 6: A comparison of solutions in three dimensions when $\lambda=10^{4}$: the exact solution (a), the solution computed by the standard PINN (b), and the solution obtained by our method (c).
@@
Figure 7: A comparison of learning curves with respect to the relative $L^{2}$-error between the standard PINN and our method for various values of $\lambda$ in three dimensions.
@@ 4.2 Example 2 (Three-dimensional case)
where $\boldsymbol{x}=(x,y,z)\in\Omega$. As before, we utilize a neural network with four hidden layers. Each layer consists of 128 neurons, and the GELU function is used as the activation function. To compute the solution, we aim to minimize the proposed loss function (3.5) with coefficients $\delta_{r}=0.05$, $\delta_{s}=1$ and $\delta_{t}=300$, and numbers of random samples $N_{r}=N_{s}=15625$ and $N_{t}=1536$, and once again we adopt the Adam optimizer.
As in the two-dimensional case, Figure 6 presents a comparison between the solutions predicted by the standard PINN and those obtained using our proposed method when $\lambda=10^{4}$. As depicted in the figure, the standard PINN produces a solution that is nearly zero with a $6.95$e11-1- 1 relative $L^{2}$-error, indicating its failure to accurately approximate the solution of the given problem. On the other hand, our proposed method yields solutions that closely match the exact solution, achieving the $9.85$e33-3- 3 relative $L^{2}$ error. Figure 7 shows the comparison of learning curves for both the standard PINN and our method for different values of $\lambda$. As before, the standard PINN fails to accurately compute the solution as $\lambda$ increases, whereas our method consistently generates accurate numerical solutions. These experiments indicate that, particularly in the nearly incompressible case, the decomposition-based approach significantly improves the solution accuracy compared to the standard PINN, and the proposed method is robust against the value of $\lambda$ also in a three-dimensional domain.
@@
Figure 8: A comparison of solutions for the case of variable Lamé parameters with $\Lambda=10^{4}$: the exact solution (a), the solution computed by the standard PINN (b), and the solution obtained by our method (c).
@@
Figure 9: A comparison of learning curves with respect to the relative $L^{2}$-error between the standard PINN and our method for the case of variable Lamé parameters with various values of $\Lambda$.
@@ 4.3 Example 3 (Inhomogeneous material)
For this experiment, we consider a neural network architecture composed of four hidden layers, each consisting of 64 neurons, with GELU activation functions. The training procedure employs the decomposed loss function (3.5), with weighting parameters set as $\delta_{r}=0.05$, $\delta_{s}=1$ and $\delta_{t}=20$, along with sampling sizes $N_{r}=5000$, $N_{s}=5000$, and $N_{t}=400$, respectively. The optimization procedure is carried out using the Adam optimizer.
Figure 8 compares the predicted solutions obtained by the standard PINN and the proposed method for the current variable Lamé parameter scenario for $\Lambda=10^{4}$. It can be clearly observed from the figure that the standard PINN produces solutions very close to zero, yielding a relative $L^{2}$-error of $9.95$e11-1- 1, thus failing to predict the solution accurately. In contrast, our proposed method provides a significantly improved prediction closely matching the exact solution, with a relative $L^{2}$-error reduced to $8.64$e33-3- 3. To further evaluate the effectiveness of the proposed approach across different material parameters, we compared learning curves for varying values of $\Lambda$, computing the relative $L^{2}$-error at each epoch. These comparisons are summarized in Figure 9. As illustrated in Figure 9 (a), when $\Lambda=10$, indicating negligible locking effects, the training performances of the standard PINN and our proposed approach are comparable. However, as $\Lambda$ increases, as shown in Figures 9 (b) and (c), the proposed method exhibits distinctly superior performance in terms of accuracy. Notably, Figure 9 (c) (when $\Lambda=10^{4}$, the inhomogeneous material becomes nearly incompressible) highlights that the standard PINN fails to converge, whereas the proposed approach maintains effective training and generates consistently accurate predictions.
@@ 4.4 Example 4 (Parametric problem)
In the practical application of elasticity, both the Young’s modulus $E$ and the Poisson’s ratio $\nu$ are typically obtained from experimental measurements of the material being modeled. Owing to this, it is natural to express the Lamé coefficients $\lambda$ and $\mu$ in terms of $E$ and $\nu$ via the following relation:
@@
and train a model providing rapid solution predictions corresponding to varying values of $\mu$ and $\lambda$. In other words, the aim in this example is to look for a parametric solution ${\bf u}$ as a function of $E$ and $\nu$ as well as the physical variable $\boldsymbol{x}\in\Omega$. In this formulation, the parameters $\boldsymbol{p}=(E,\nu)$ are sampled from the parametric domain $\mathcal{P}=(2,4)\times(0.1,0.5)$, which encompasses, in particular, the near-incompressibility regime.
For the experiment, we employ the same setting as used in Example 1, where we consider the linear elasticity equation (1.1) with the true solution
${\bf u}(\boldsymbol{x})=\begin{bmatrix}u_{1}(x,y)\\ u_{2}(x,y)\end{bmatrix}=\begin{bmatrix}\bigl{(}\cos(2x)-1\bigr{)}\sin(2y)\\ \bigl{(}1-\cos(2y)\bigr{)}\sin(2x)\end{bmatrix}+\frac{\sin(x)\sin(y)}{\lambda}\begin{bmatrix}1\\ 1\end{bmatrix},$
for ${\boldsymbol{x}}=(x,y)\in\Omega=(0,\pi)^{2}$. Utilizing the notation established in Section 2.2, we choose $\delta_{r}=1$ and $\delta_{b}=20$ in (2.5), that is, the parametric loss function $\mathcal{L}_{\rm para}=\mathcal{L}_{{\rm{para}},\,r}+20\mathcal{L}_{{\rm{para}},\,b},$ where $\mathcal{L}_{{\rm{para}},\,r}$ and $\mathcal{L}_{{\rm{para}},\,b}$ are defined in (2.4). The neural network architecture consists of four hidden layers with 64 neurons per layer and employs GELU as an activation function. For the training procedure, we used the Adam optimizer and generated sampling points from a uniform distribution, with $N_{pb}\times N_{b}=30{,}000$ and $N_{pr}\times N_{r}=100{,}000$ used for the boundary and residual components, respectively.
@@
Figure 10: Error surfaces for the relative $L^{2}$-errors in the parametric solution ${\bf u}(\cdot;E,\nu)$ across varying values of $(E,\nu)$.
To further demonstrate the effectiveness of the proposed method in mitigating the locking phenomenon, we evaluated the error with respect to varying values of $\nu$ while keeping $E$ fixed. Specifically, we divided the range of $\nu$ into seven intervals: $(0.1,0.2)$, $(0.2,0.3)$, $(0.3,0.4)$, $(0.4,0.45)$, $(0.45,0.49)$, $(0.49,0.499)$ and $(0.499,0.4999)$. For each interval, 100 random values of $\nu$ were sampled, and the averaged relative $L^{2}$-error was measured. As shown in the Figure 11, our method consistently produced accurate solutions across all intervals. In particular, even when the $\nu$ is close to $0.5$ where locking typically occurs, the proposed method produces accurate and stable solution predictions.
Figure 11: Comparison of the average relative $L^{2}$-errors in the parametric solution ${\bf u}(\cdot;E,\nu)$ of the standard PINN and our method over $100$ samples in each interval with different range of $\nu$.
@@ 5 Conclusion
In this paper, we investigate the linear elasticity equations within the nearly incompressible regime and propose a novel and robust approach utilizing PINNs. Unlike the conforming FEM, where locking (non-robustness) phenomena occur primarily due to approximation limitations, locking in PINNs arises predominantly from a pronounced imbalance between the Lamé parameters, thereby causing significant difficulties in optimization. To resolve this issue, we introduce a decomposition of the governing equations into two carefully structured subsystems, separating the effect of each parameter. Our proposed approach concurrently addresses forward and inverse problems during the training, automatically identifying external forces and boundary conditions suitable for the considered scenarios. The efficiency and robustness of our methodology are validated through comprehensive numerical experiments, including both two- and three-dimensional examples, problems involving variable coefficients, and parametric settings. The results consistently demonstrate the capability of our method to effectively handle elastic materials that become nearly incompressible, highlighting its potential applicability to complex elasticity problems in computational mechanics.
A rigorous convergence analysis to establish the reliability of the method and an application of our method to more practical problems in real-world scenarios would form an interesting future research direction. Moreover, the decomposition-based approach we propose is expected to be applicable to a broader class of problems in which the quality of numerical approximation is highly sensitive to parameter values, such as singular perturbation problems involving small diffusion parameters, or the Navier–Stokes equations with large Reynolds numbers. These possibilities will be explored in future works.

### v2

@@ 6 Numerical Experiments
Figure 5: A comparison of solutions in two dimensions when $\lambda=10^{5}$: the exact solution (a), the solution computed by the standard PINN (b), and the solution obtained by our method (c). The experiment shows that standard PINNs fail to approximate the true solution in this example.
@@
Figure 6: A comparison of learning curves with respect to the relative $L^{2}$-error between the standard PINN and our method for various values of $\lambda$ in two dimensions. The experiment shows that for $\lambda=10^{5}$, the standard PINN approach does not yield an approximation to the true solution in this example.
In this section, we conduct a series of numerical experiments to assess the effectiveness of the proposed method. We compare a model trained with the standard PINN against a model trained using the proposed decomposition-based approach. To provide a comprehensive evaluation of the proposed method when the Lamé coefficients $\mu$ and $\lambda$ in the elasticity equations are constant, including scenarios with large $\lambda$ (nearly incompressible material), we perform some numerical experiments in both two and three dimensions. We then investigate the case of variable $\lambda$ and $\mu$ (inhomogeneous material). Finally, we will also conduct experiments on a parametric PINN, which enables the rapid prediction of solutions for varying Lamé coefficients without the need for retraining. To ensure a fair comparison, all other settings, including the model architecture, initialization, and optimization algorithm, remain unchanged.
@@ 6.1 Example 1 (Two-dimensional case)
${\bf u}(\boldsymbol{x})=\begin{bmatrix}u_{1}(x,y)\\ u_{2}(x,y)\end{bmatrix}=\begin{bmatrix}\bigl(\cos(2x)-1\bigr)\sin(2y)\\ \bigl(1-\cos(2y)\bigr)\sin(2x)\end{bmatrix}+\frac{\sin(x)\sin(y)}{\lambda}\begin{bmatrix}1\\ 1\end{bmatrix}$, (6.1)
where ${\boldsymbol{x}}=(x,y)\in\Omega$. For a baseline model, we implement a neural network consisting of four hidden layers, each containing 50 neurons, activated by the GELU function. For training, as described before, we adopt the decomposed loss function (4.5) with weights $\delta_{r}=0.05$, $\delta_{s}=1$ and $\delta_{b}=20$, and number of sampling points $N_{r}=5000$, $N_{s}=5000$ and $N_{b}=400$. We use the Adam optimizer as an optimization algorithm.
Figure 5 illustrates the solutions predicted by a standard PINN and our proposed method when $\lambda=10^{5}$ (nearly incompressible material). As shown in the figure, the solution obtained using the original PINN is close to zero with 1.011.01e0 relative $L^{2}$-error, failing to accurately predict the solution of the given problem. In contrast, the solution computed using our proposed method closely approximates the exact solution with 1.651.65e3-3 relative $L^{2}$-error. To compare the learning curves for different values of $\lambda$, we conducted training using both the standard PINN and our proposed method for various values of $\lambda$, computing the relative $L^{2}$-error at each training epoch. The results of this experiment are presented in Figure 6. As observed from Figure 6 (a), when $\lambda=10$ (no locking is anticipated), there is no significant difference in training efficiency between the original PINN and our proposed method. However, as $\lambda$ is getting large, it is clear from (b) and (c) of Figures 6 that our method demonstrates superior training efficiency and accuracy compared to the standard PINN. It is noteworthy that, as we can see from Figure 6 (c) (when $\lambda=10^{5}$, the material becomes nearly incompressible), the standard PINN fails to train due to the non-robustness phenomenon in the nearly incompressible case, whereas our proposed method continues to achieve effective training with high accuracy.
@@
Figure 7: A comparison of solutions in three dimensions when $\lambda=10^{4}$: the exact solution (a), the solution computed by the standard PINN (b), and the solution obtained by our method (c). The experiment shows that standard PINNs fail to approximate the true solution in this example.
@@
Figure 8: A comparison of learning curves with respect to the relative $L^{2}$-error between the standard PINN and our method for various values of $\lambda$ in three dimensions. The experiment shows that for $\lambda=10^{4}$, the standard PINN approach does not yield an approximation to the true solution in this example.
@@ 6.2 Example 2 (Three-dimensional case)
where $\boldsymbol{x}=(x,y,z)\in\Omega$. As before, we utilize a neural network with four hidden layers. Each layer consists of 128 neurons, and the GELU function is used as the activation function. To compute the solution, we aim to minimize the proposed loss function (4.5) with coefficients $\delta_{r}=0.05$, $\delta_{s}=1$, and $\delta_{b}=300$, and numbers of random samples $N_{r}=N_{s}=15625$ and $N_{b}=1536$, and once again we adopt the Adam optimizer.
As in the two-dimensional case, Figure 7 presents a comparison between the solutions predicted by the standard PINN and those obtained using our proposed method when $\lambda=10^{4}$. As depicted in the figure, the standard PINN produces a solution that is nearly zero with a 6.956.95e1-1 relative $L^{2}$-error, indicating its failure to accurately approximate the solution of the given problem. On the other hand, our proposed method yields solutions that closely match the exact solution, achieving the 9.859.85e3-3 relative $L^{2}$ error. Figure 8 shows the comparison of learning curves for both the standard PINN and our method for different values of $\lambda$. As before, the standard PINN fails to accurately compute the solution as $\lambda$ increases, whereas our method consistently generates accurate numerical solutions. These experiments indicate that, particularly in the nearly incompressible case, the decomposition-based approach significantly improves the solution accuracy compared to the standard PINN, and the proposed method is robust against the value of $\lambda$ also in a three-dimensional domain.
@@
Figure 9: A comparison of solutions for the case of variable Lamé coefficients with $\Lambda=10^{4}$: the exact solution (a), the solution computed by the standard PINN (b), and the solution obtained by our method (c). The experiment shows that standard PINNs fail to approximate the true solution in this example.
@@
Figure 10: A comparison of learning curves with respect to the relative $L^{2}$-error between the standard PINN and our method for the case of variable Lamé coefficients with various values of $\Lambda$. The experiment shows that for $\lambda=10^{4}$, the standard PINN approach does not yield an approximation to the true solution in this example.
@@ 6.3 Example 3 (Inhomogeneous material)
For this experiment, we consider a neural network architecture composed of four hidden layers, each consisting of 64 neurons, with GELU activation functions. The training procedure employs the decomposed loss function (4.5), with weighting parameters set as $\delta_{r}=0.05$, $\delta_{s}=1$ and $\delta_{b}=20$, along with sampling sizes $N_{r}=5000$, $N_{s}=5000$, and $N_{b}=400$, respectively. The optimization procedure is carried out using the Adam optimizer.
Figure 9 compares the predicted solutions obtained by the standard PINN and the proposed method for $\Lambda=10^{4}$. It can be clearly observed from the figure that the standard PINN produces solutions very close to zero, yielding a relative $L^{2}$-error of 9.959.95e1-1, thus failing to predict the solution accurately. In contrast, our proposed method provides a significantly improved prediction closely matching the exact solution, with a relative $L^{2}$-error reduced to 8.648.64e3-3. To further evaluate the effectiveness of the proposed approach across different material parameters, we compared learning curves for varying values of $\Lambda$, computing the relative $L^{2}$-error at each epoch. These comparisons are summarized in Figure 10. As illustrated in Figure 10 (a), when $\Lambda=10$, indicating negligible locking effects, the training performances of the standard PINN and our proposed approach are comparable. However, as $\Lambda$ increases, as shown in Figures 10 (b) and (c), the proposed method exhibits distinctly superior performance in terms of accuracy. Notably, Figure 10 (c) (when $\Lambda=10^{4}$, the inhomogeneous material becomes nearly incompressible) highlights that the standard PINN fails to converge, whereas the proposed approach maintains effective training and generates consistently accurate predictions.
@@ 6.4 Example 4 (Parametric problem)
In the practical application of elasticity, both the Young’s modulus $E$ and the Poisson’s ratio $\nu$ are typically obtained from experimental measurements of the material being modeled. Owing to this, it is natural to express $\lambda$ and $\mu$ in terms of $E$ and $\nu$ via the following relation:
@@
and train a model providing rapid solution predictions corresponding to varying values of $\mu$ and $\lambda$. In other words, the aim in this example is to look for a parametric solution $\bf u$} as a function of $E$ and $\nu$. In this formulation, the parameters $\boldsymbol{p}=(E,\nu)$ are sampled from the parametric domain $\mathcal{P}=(2,4)\times(0.1,0.5)$, which encompasses, in particular, the near-incompressibility regime.
For the experiment, we employ the same setting as used in Example 1, where we consider the linear elasticity equation (1.1) with the true solution given in (6.1). Utilizing the notation established in Section 2.2, we choose $\delta_{r}=1$ and $\delta_{b}=20$ in (2.5), that is, the parametric loss function $\mathcal{L}_{\rm para}=\mathcal{L}_{{\rm{para}},\,r}+20\mathcal{L}_{{\rm{para}},\,b},$ where $\mathcal{L}_{{\rm{para}},\,r}$ and $\mathcal{L}_{{\rm{para}},\,b}$ are defined in (2.4). The neural network architecture consists of four hidden layers with 64 neurons per layer and employs GELU as an activation function. For the training procedure, we used the Adam optimizer and generated sampling points from a uniform distribution, with $N_{pb}\times N_{b}=30{,}000$ and $N_{pr}\times N_{r}=100{,}000$ used for the boundary and residual components, respectively.
@@
Figure 11: Error surfaces for the relative $L^{2}$-errors in the parametric solution ${\bf u}(\cdot;E,\nu)$ across varying values of $(E,\nu)$. The experiment shows that for standard PINN, there is a sharp increase in the error as the material becomes nearly incompressible (i.e., $\nu$ approaches $1/2$).
To further demonstrate the effectiveness of the proposed method in mitigating the locking phenomenon, we evaluated the error with respect to varying values of $\nu$ while keeping $E$ fixed. Specifically, we divided the range of $\nu$ into seven intervals: $(0.1,0.2)$, $(0.2,0.3)$, $(0.3,0.4)$, $(0.4,0.45)$, $(0.45,0.49)$, $(0.49,0.499)$ and $(0.499,0.4999)$. For each interval, 100 random values of $\nu$ were sampled, and the averaged relative $L^{2}$-error was measured. As shown in the Figure 12, our method consistently produced accurate solutions across all intervals. In particular, even when the $\nu$ is close to $0.5,$ where locking typically occurs, the proposed method produces accurate and stable solution predictions.
Figure 12: Comparison of the average relative $L^{2}$-errors in the parametric solution ${\bf u}(\cdot;E,\nu)$ of the standard PINN and our method over $100$ samples in each interval with different range of $\nu$. The experiment shows that a standard PINN approach is sensitive to the value of $\nu$, especially in the nearly incompressible regime (i.e., $\nu$ approaches $1/2$).
@@ 7 Conclusion
In this paper, we investigate the linear elasticity equations within the nearly incompressible regime and propose a novel and robust approach utilizing PINNs. Unlike the conforming FEM, where locking (non-robustness) phenomena occur primarily due to approximation limitations, locking in PINNs arises predominantly from a pronounced imbalance between the Lamé coefficients, thereby causing significant difficulties in optimization. To resolve this issue, we introduce a decomposition of the governing equations into two carefully structured subsystems, separating the effect of each parameter. Our proposed approach concurrently addresses forward and inverse problems during the training, automatically identifying external forces and boundary conditions suitable for the considered scenarios. The efficiency and robustness of our methodology are validated through comprehensive numerical experiments, including both two- and three-dimensional examples, problems involving variable coefficients, and parametric settings. The results consistently demonstrate the capability of our method to effectively handle elastic materials that become nearly incompressible, highlighting its potential applicability to complex elasticity problems in computational mechanics.
An application of our method to more practical problems in real-world scenarios would form an interesting future research direction. Moreover, the decomposition-based approach we propose is expected to be applicable to a broader class of problems in which the quality of numerical approximation is highly sensitive to parameter values, such as singular perturbation problems involving small diffusion parameters, or the Navier–Stokes equations with large Reynolds numbers. These possibilities will be explored in future works.
Acknowledgments
Seungchan Ko and Sanghyeon Park are supported by National Research Foundation of Korea Grant funded by the Korean Government (RS-2023-00212227). Josef Dick, Quoc Thong Le Gia, and Kassem Mustapha are partially supported by the ARC Grant DP220101811.